OpenStudy (marigirl):

An infestation of grubs in a farmer's paddock has been increasing since he purchased the property. After 3 years of owning the property, the infestation of grubs reached 60 grubs per m^2. A survey done 5 years later (t=8) indicated there were now 105 grubs per m^2. The equation that models the infestation can be represented by I=Ak^t. Find the constants A and k. please help

OpenStudy (marigirl):

(3, 60) (8, 105)

OpenStudy (daniellelovee):

105/60=Ak8/Ak3 7/4=k8-3 7/4=k5 k=?

OpenStudy (marigirl):

my k value is 1.118 (which is the answer in the back of the book). My A value is 42.88 Where as the back of the book says 42.94.. is my answer ok?

OpenStudy (danjs):

hmm,

OpenStudy (danjs):

\[\large 60=A*k^3~~~~~and~~~~~105=A*k^8\]

OpenStudy (daniellelovee):

looks correct to me :)

OpenStudy (daniellelovee):

and for A I got A=42.887 which I guess got simplified and that is why your book says 4.92

OpenStudy (danjs):

Solve second for A, A=105/k^8, and put in first \[60=\frac{ 105 }{ k^5 }~~~~~k=(\frac{ 105 }{ 60 })^{1/5} \approx 1.118\]

OpenStudy (danjs):

It depends on how many decimals you use for the k value

OpenStudy (daniellelovee):

yes...but why are you repeating the steps that we have said?

OpenStudy (danjs):

exact numbers \[\large k=\frac{ 7^{1/5}*2^{3/5} }{ 2 }\] \[\large A=\frac{ 60*7^{2/5}*2^{6/5} }{ 7 }\]

OpenStudy (danjs):

That has A=42.88727... , just depends on if you round off k